A. The following is a macroeconomic model of Utopia. The Utopian currency is the Utopian dollar (#).
Y = C + I + G + NX (National Income Identity)
C = C0 + cYd
(Consumption Function)
I = I0 (Investment Function)
G = G0 (Government Expenditure Function)
X = 20 (Export Expenditure)
M = 35 (Import Expenditure)
Y
d
= Y – T (Disposable Income)
Required:
i. Compute the value of the balance of payments (BOP). (1 mark)
ii. State the BOP position. (1 mark)
iii. Explain the BOP position you found in (ii) (1 mark)
iv. State and explain one (1) strategy for the Utopian government to improve the BOP position. (2 marks)
1
Expert's answer
2020-05-28T11:40:20-0400

i) "NX=X-N=20-35=-15"

NX-BOP

ii) The BOP is negative and reduces GDP.

iii) This situation shows the Utopia buys more than sells and the goverment depends on other countries.

iv) Goverment should subsidies for sectors producing goods or services that account for most of the imports.